4.Roulette. A roulette wheel has 38 equal size slots that a small ball can settle in after rolling around the wheel. The slots are labeled 0, 00, and 1 through 36.

a.What is the probability of the ball landing on 0 or 00?

b.The 0 and 00 slots are colored green. Of the other 36 slots, 18 are colored red and 18 black. What is the probability of the ball landing on black?
Answer #1

probability = number of favorable cases/total number of cases

a. the probability of the ball landing on 0 or 00 = 2/38 = 0.05263157894

b. the probability of the ball landing on black = 18/38 = 0.47368421052

press like